OpenStudy (anonymous):

Simplify cos2 Θ(1 + tan2 Θ)

OpenStudy (anonymous):

I'm supposed to be using the Pythagorean Identity to do this...

OpenStudy (anonymous):

@Abhisar

OpenStudy (science0229):

Does the '2' represent square? Like \[\cos^2(1+\tan^2)\]

OpenStudy (anonymous):

Yes, it does. Sorry about that!

OpenStudy (anonymous):

You got the format correct the second time around.

OpenStudy (anonymous):

Well, your first.

OpenStudy (science0229):

Do you know any trig pythagorean identities?

OpenStudy (anonymous):

Ummmm, sin^2+cos^2=1

OpenStudy (anonymous):

I think that's how it goes

OpenStudy (science0229):

With that, this problem can be solved!

OpenStudy (science0229):

You just have to change it around a bit

OpenStudy (bradely):

cos^2 Θ(1 + tan^2 Θ) cos^2 Θ(1 +( sin^2 Θ/cos^2 Θ)) cos^2 Θ((cos^2 Θ + sin^2 Θ)/cos^2 Θ)) =1 source: http://www.mathskey.com/question2answer/

OpenStudy (science0229):

Look at this and see if you understand:\[\frac{ \sin^2+\cos^2 }{ \cos^2 }=\frac{ 1 }{ \cos^2 }\]

OpenStudy (science0229):

\[\frac{ \sin^2+\cos^2 }{ \cos^2 }=\frac{ \sin^2 }{ \cos^2 }+\frac{ \cos^2 }{ \cos^2 }=\]

OpenStudy (anonymous):

Couldn't you simplify that to tan^2+cos=1/cos^2?

OpenStudy (science0229):

Exactly!

OpenStudy (science0229):

Now substitute that into the original equation to get...

OpenStudy (anonymous):

Yes! I love it when shot's in the dark hit the target

OpenStudy (anonymous):

Whoa, why are we substituting? I thought we were simplifying the same equation the whole time.

OpenStudy (anonymous):

I'm sorry :/

OpenStudy (science0229):

Wait...

OpenStudy (anonymous):

Ok...

OpenStudy (science0229):

I almost missed that! \[\frac{ \sin^2 }{ \cos^2 }+\frac{ \cos^2 }{ \cos^2 }=\tan^2+1\]

OpenStudy (science0229):

which equals to 1/cos^2

OpenStudy (anonymous):

Thats saying is that Tan^2=1

OpenStudy (anonymous):

O crud, forgot a one

OpenStudy (anonymous):

Nevermind

OpenStudy (science0229):

Nope. \[\tan^2+1=\]

OpenStudy (anonymous):

Isn't it saying 1=1?

OpenStudy (science0229):

\[\frac{ 1 }{ \cos^2 }\]

OpenStudy (bradely):

u can write 1+tan^2 (theta) = sec^2(theta)

OpenStudy (anonymous):

Please do

OpenStudy (bradely):

sec^2(theta) will be write 1/cos^2(theta)

OpenStudy (bradely):

cos^2(theta)*(1/cos^2(theta)) =1

OpenStudy (anonymous):

I'm hopelessly confused. I'm so sorry

OpenStudy (anonymous):

Do we still work at \(cos^2 (1+ tan^2)\)??

OpenStudy (anonymous):

Yes, we are

OpenStudy (anonymous):

We're trying to simplify it using Pythagorean Identity/ies

OpenStudy (anonymous):

ok, distribute cos^2 into the bracket, we have \(cos^2 + cos^2* tan^2\) ok then?

OpenStudy (anonymous):

now, tan =\(\dfrac{sin}{cos}\) so that \(tan^2 =\dfrac{sin^2}{cos^2} \)

OpenStudy (anonymous):

Why are there two cos^2 in your first reply?

OpenStudy (anonymous):

replace to expression, you have \(cos^2 + cos^2*\dfrac{sin^2}{cos^2}= cos^2 + sin^2 =1\)

OpenStudy (anonymous):

I'm sorry, I'm taking an online course that never explained any of this, and math isn't my strong suit anyways.

OpenStudy (anonymous):